I would be truly grateful if a tutor could show me how to do these prblems or at least give me the equation to set them up.
With probability 0.05 a robotic reporter that works for MEOW NEWS will make a false statement.
Suppose the reporter produced 1900 statements last week.
12. (1 point) What is the expected number of false statements, and the standard deviation?
13. (1 point) What is the probability that 85 or fewer of those statements were false? Round answer to
four places after the decimal point.
14. (1 point) What is the probability that 124 or more of those statements were false? Round answer to
four places after the decimal point.
-------------------------------------------------------------------

Answer by stanbon(75887)

With probability 0.05 a robotic reporter that works for MEOW NEWS will make a false statement.
Suppose the reporter produced 1900 statements last week.
12. (1 point) What is the expected number of false statements, and the standard deviation?
---
E(x) = np = 1900*0.05 = 5*19 = 95
std = sqrt(npq) = sqrt[95*0.95] = 9.5
-----------------------------------
13. (1 point) What is the probability that 85 or fewer of those statements were false? Round answer to four places after the decimal point.
---
P(0<= x <= 85) = binomcdf(1900,0.05,85) = 0.1586
-----------------------------------------------------
14. (1 point) What is the probability that 124 or more of those statements were false? Round answer to four places after the decimal point.
P(124 <= x <= 1900) = 1 - binomcdf(1900,0.05,123) = 0.0019
==============================================================
Cheers,
Stan H.
-------------------------------------------------------------------